For the following function C computes the cost in millions of dollars of implementing a city recycling project when x percent of the citizens participate C(x)=1.3x/100-x. Using this model approximatelywhat is the cost if 40% of the citizens participate? Include units with your answer: For the following function C computes the cost in millions of dollars of implementing a city recycling project when x percent of the citizens participate C(x)=1.3x/100-x. Using this model approximatelywhat is the cost if 40% of the citizens participate? Include units with your answer
Answer by stanbon(18991)

For the following function C computes the cost in millions of dollars of implementing a city recycling project when x percent of the citizens participate C(x)=1.3x/100-x. Using this model approximatelywhat is the cost if 40% of the citizens participate? Include units with your answer
-------------------------
C(40) = 1.3*40/(100-40)
C(40) = 52/60 = 13/15 = 0.866666.. million dollars
= $866,666.67
================
Cheers,
Stan H.